At the beginning of 2019, Robotics Inc. acquired a manufacturing facility for $13.8 million. $10.8 million of the purchase price was allocated to the building. Depreciation for 2019 and 2020 was calculated using the straight-line method, a 20-year useful life, and a $2.8 million residual value. In 2021, the company switched to the double-declining-balance depreciation method.

What is depreciation on the building for 2021? (Do not round intermediate calculations. Round answer to the nearest whole dollar.)
Answer #1

Answer:

A) Depreciation as per SLM = (Cost -Salvage)/Useful Life = (1,08,00,000 - 28,00,000)/20 = $320,000
Therefore Depreciation for 2019 =$400,000
Depreciation for 2020 =$400,000
Accumulated Depreciation =$800,000.
Now, Depreciation method changes to DDM
Undepreciated cost =108,00,000-800,000=$1,00,00,000
Residual Value = $28,00,000
Remaining life = 18 years
Declining Rate = (1/18)*2 = 11.1111111%
Revised annual Depreciation = 11.11111111%*1,00,00,000 = $11,11,111.